OpenStudy (anonymous):

please help find the solution set for 3+2(x-4)=5-2(3-2x)

OpenStudy (anonymous):

First you distribute 3+2x-8=5-6+4x Then combine like terms 2x-5=-1+4x Then you get x by itself by minusing 2x from one side. I prefer to keep things positive so I'll minus 2 from 4 -5=-1+2x The add 1 to both sides -4=2x Dive by 2 on each side -2=x

OpenStudy (anonymous):

Divide, not dive, sprry

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome